OpenStudy (anonymous):

What is the surface area of a cube with side length 9 kilometers? two hundred sixteen square kilometers three hundred twenty-four square kilometers four hundred eighty-six square kilometers seven hundred twenty-nine square kilometers @matricked @dg2 @xanny6

OpenStudy (anonymous):

surface area of the cube=6a^2 a=9 plug in the value in the formula

OpenStudy (anonymous):

use surface area = 6*a^2 where a=length of side

OpenStudy (anonymous):

here a=9 so u may get four hundred eighty-six square kilometers